OpenStudy (anonymous):

@HELP_ME! f(x)=5/x^3 find the antiderivative

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{5}{x^3}\]If we use an exponent rule, we can write our function like this,\[\Large\rm f(x)=5x^{-3}\]This allows us to more easily apply our power rule.

zepdrix (zepdrix):

Errr.. power rule for integration, whatever you want to call it :)

OpenStudy (anonymous):

ok so far so good :)

zepdrix (zepdrix):

So we're using the fact that the antiderivative of \(\Large\rm x^n\) is \(\Large\rm \frac{1}{n+1}x^{n+1}\)

jimthompson5910 (jim_thompson5910):

use \[\Large \int x^ndx = \frac{1}{n+1}x^{n+1}+C\]

zepdrix (zepdrix):

Power increases by 1, then divide by that new power.

OpenStudy (anonymous):

so when the power increases in become -4 right?

zepdrix (zepdrix):

Ah! So it's easy to get tripped up when we're dealing with negative numbers. So our power will become -3+1

OpenStudy (anonymous):

ok so its -2 then

zepdrix (zepdrix):

Mmm looks good

OpenStudy (anonymous):

so.. 5x^-2/-2 +c?

zepdrix (zepdrix):

Good good good. It might be a good idea to put it back into the form that it was given to you in:\[\Large\rm -\frac{5}{2x^{2}}+c\]

jimthompson5910 (jim_thompson5910):

correct, \[\Large -\frac{5}{2}x^{-2} + C\]

zepdrix (zepdrix):

Not a big deal though :)

OpenStudy (anonymous):

@zepdrix yea my textbook answer puts it back into the form it was given but im glad its still correct lol

OpenStudy (anonymous):

thank you both

zepdrix (zepdrix):

np